Joey’s is a manufacturer with annual sales of 5000 units. It costs them $2 per unit and they sell at a price of $6 each. Fixed costs to run the business are 160 000 per year.

  1. What is the current total contribution?

Joey’s is considering increasing the price to $7.50/unit and increasing its advertising to $4000. This would result in a sales increase of 33%.

  1. What is the total new contribution?

  

  1. What is incremental contribution?


  1. Should you go ahead with the adjustment?
Answer #1

Current Total Contribution = (Selling Priceper unit-Variable Cost per unit)*Sales Quantity = (6-2)*5000 = $20000

New Sales Quantity = 5000*(1+0.33) = 6650

Total New Contribution = (New Selling Price-New Variable Cost)*New Sales Quantity = (7.5-2)*6650 = $36575

Incremental Contribution = New Contibution-Current Contribution = 36575-20000 = $16575

Additional PROFIT due to adjustment = Incremental Contribution-Additional Advertisement = 16575-4000 = $12575

Additonal Profit of $12575 is achieved due to the adjustment. Therefore, Adjustment SHOULD BE made.